A 515kg roller coaster is at the bottom of a loop with a radius of 10m. If the speed at the bottom of the loop is 20m/s, what is the force of the track pushing up on the vehicle at this point?



1
Expert's answer
2021-02-01T09:37:33-0500

v = 20 m/s

R = 10 m

m = 515 kg

"N - mg = ma_c \\

a_c = \frac{v^2}{R} \\

N = m(\frac{v^2}{R} + g) \\

= 515(\frac{400}{10}+10) \\

= 515 \times 50 \\

= 25750 \;N"

Answer: 25750 N